The length of a rectangle is two feet more than four times its width. find the width given the area is 42 feet.
what i have done:
42=4(x+2)
42=4x+8
0=4x-34

Answer by oscargut(2103)

width=x
length=4x+2
x(4x+2)=42
4x^2+2x-42=0
2x^2+x-21=0
solutions are 3 and -7/2
so x=3
and 4x+2=14
Answer: width is 3 feets